# Time-Series Foundation Models for ISP Traffic Forecasting

### v1

@@ I Introduction
  • RQ2: Cross-level generalization. Can TTM generalize effectively to ISP traffic across hierarchical levels (institutions, subnets, and IP addresses), without retraining?
  • RQ3: Horizon scaling. How does performance evolve as the forecasting horizon increases (from hours to days ahead)?
@@
  • Evaluation of TTM on real ISP traffic. We apply IBM’s TTM to CESNET-TimeSeries24 and evaluate its performance under the benchmark protocol.
@@
  • Negative result on few-shot learning. We systematically evaluate few-shot fine-tuning and exogenous features, and find negligible improvement, highlighting the robustness of TTM’s pretrained representations for ISP traffic.
@@ II-A Statistical Forecasting
Early network-traffic forecasting methods relied on statistical time-series models such as exponential smoothing and the ARIMA models [13]. These approaches remain appealing for their interpretability and low computational cost, yet a linear correlation structure is assumed among the time series values, and therefore, no nonlinear patterns can be captured by the ARIMA model. The approximation of linear models to complex real-world problems is not always satisfactory. And many researchers have argued that real-world systems are often nonlinear [14]. Hybrid extensions combining ARIMA with neural models such as ANN and LSTM [15, 16] have been explored to capture both of linear and non-linear patterns, but their improvements often come at the expense of increased model complexity and extensive parameter tuning.
@@ II-E Gap in ISP Forecasting
While foundation models have achieved promising results in finance, energy, and transportation [30], their application to Internet traffic remains largely unexplored. [1] identifies generalization and transfer learning as open challenges in network-traffic prediction.
@@
Our study addresses these questions by systematically evaluating TTM on the CESNET-TimeSeries24 benchmark, assessing zero-shot, few-shot, and exogenous-feature settings to determine whether pretrained forecasting models can match or surpass specialized deep-learning baselines while maintaining low computational overhead.
@@ III Methodology
This section describes the experimental pipeline used to evaluate TTM-R2 (where R2 denotes the model’s release version; TTM is used throughout the paper) for ISP traffic forecasting. As illustrated in Figure 1, the workflow comprises five stages: (1) dataset preprocessing and window generation, (2) model configuration and context–horizon setup, (3) zero-shot and few-shot execution, (4) metric computation across overlapping windows, and (5) efficiency benchmarking on CPU-only mode. This design ensures reproducible, temporally consistent evaluation across both time resolutions and hierarchical aggregation levels.
@@ III-A Dataset and Forecasting Targets
Each record captures multiple traffic attributes including n_flows, n_packets, n_bytes, tcp_udp_ratio_bytes, and avg_duration etc.. Data are collected at a base granularity of 10 minutes and further aggregated to hourly resolution by temporal averaging, resulting in two temporal granularities:
@@
Among the available metrics, we focus primarily on n_bytes, representing the total traffic volume per interval. We initially experimented with multiple target attributes and observed consistent performance trends across them. We therefore report n_bytes as the representative target for clarity and comparability. This metric directly reflects bandwidth utilization and network load, making it operationally relevant for applications such as capacity planning, congestion management, and anomaly detection in ISP networks.
@@ III-F Exogenous Variables
We test with/without exogenous signals (binary weekend and holiday flags). Default is off (univariate).
@@ IV Results and Discussion
IV-A Effect of Context Length
@@
For hourly forecasts ($H=96$), the RMSE remains nearly unchanged as $L$ increases from 512 to 1536, indicating that the model already encodes sufficient temporal memory with moderate context. Similarly, at 10-minute resolution ($H=336$), longer lookback windows show no consistent improvement in either RMSE or $R^{2}$. These results suggest that TTM’s pretrained temporal representations capture recurring diurnal and weekly cycles even when short historical windows are used. This shows that foundation models learn strong implicit periodic priors during pretraining, reducing dependence on extensive local history.
@@
IV-B Effect of Forecast Horizon
@@
IV-C Cross-Level Generalization
Our findings with TTM align with this general trend but demonstrate greater resilience. As summarized in Table III (see Appendix), mean $R^{2}$ values remain positive and stable for institutions ($\sim$0.25) and subnets ($\sim$0.17), while declining at the IP level ($\sim$–1.39) due to extreme sparsity. Unlike the baseline study, which suggested that distinct modeling strategies may be needed at each aggregation level (e.g., recurrent models for institutions, specialized preprocessing for IPs), TTM maintains a single unified architecture across all levels, without any retraining or architectural modification.
[…]
IV-D Few-Shot Adaptation and Exogenous Variables
Figure 5 shows that as the fraction of fine-tuning data increases from 10% to 50% ($L=1024$, $H=96$), RMSE improves modestly (from $\sim$0.075 to $\sim$0.070), while $R^{2}$ remains largely stable. This indicates that limited adaptation can refine magnitude estimation but does not substantially alter the model’s overall calibration, consistent with strong pretrained temporal priors.
TABLE II: Inference and Training Efficiency Across Experimental Configurations (CPU-only, Apple M2 Pro, 16GB RAM). All values reported as mean ± standard deviation. Training time only applicable to few-shot experiments.
@@ IV-D Few-Shot Adaptation and Exogenous Variables
IV-E Deployability and Efficiency
Computational efficiency is a key requirement for operational ISP deployment. Unlike conventional deep learning models that rely on GPUs and extensive retraining, TTM delivers strong forecasting accuracy entirely on CPU hardware with minimal overhead. All experiments were executed on a MacBook Pro (M2 Pro, 16 GB RAM, CPU-only). Table II summarizes inference and training costs across configurations. For hourly forecasting at the institutional level ($L=1024$, $H=96$), TTM achieves 0.044 s per 100 points ($\sim$2,200 samples/s). Efficiency remains stable across context lengths (L=512L=512–1536) with only a 0.006 s difference.
@@ IV-F Quantitative Summary
Table III (see Appendix) reports the aggregated metrics (mean ± std across all evaluated series). TTM maintains consistent accuracy across context lengths and horizons, with RMSE around 0.025–0.056 and stable inference times (Table II) even on CPU-only mode. R² remains positive for institutional and subnet levels, with degradation at the IP level due to flat or zero-variance series. These results confirm the model’s robustness and computational efficiency across traffic hierarchies.
@@ Appendix A Additional Results
TABLE III: Forecasting Accuracy Summary Across All Experimental Configurations. Values reported as mean ± standard deviation after removing top/bottom 5% outliers.
Experiment Context L Horizon H Few-shot % RMSE
Context Length Sensitivity (Hourly, Institutions)
Hourly L=512 512 96 0 0.0566 ± 0.0269 0.2127 ± 0.1388
Hourly L=1024 1024 96 0 0.0551 ± 0.0263 0.2460 ± 0.1571
Hourly L=1536 1536 96 0 0.0551 ± 0.0265 0.2498 ± 0.1576
Cross-Level Generalization (Hourly, L=1024, H=96)
Cross: Institutions 1024 96 0 0.0551 ± 0.0263 0.2460 ± 0.1571
Cross: Subnets 1024 96 0 0.0541 ± 0.0365 0.1685 ± 0.1672
Cross: IPs 1024 96 0 0.0317 ± 0.0262 $1.3872\pm 8.6764$
Horizon Scaling (10-minute, Institutions, L=1024)
10-min H=96 1024 96 0 0.0259 ± 0.0142 0.0975 ± 0.0982
10-min H=336 1024 336 0 0.0268 ± 0.0144 0.0654 ± 0.0725
10-min H=720 1024 720 0 0.0271 ± 0.0146 0.0456 ± 0.0638
Few-Shot Learning (Hourly, Institutions, L=1024, H=96)
Few-shot: 0% (zero-shot) 1024 96 0 0.0551 ± 0.0263 0.2460 ± 0.1571
Few-shot: 10% 1024 96 10 0.0718 ± 0.0380 $0.2170\pm 0.3600$
Few-shot: 30% 1024 96 30 0.0605 ± 0.0299 0.1196 ± 0.1847
Few-shot: 50% 1024 96 50 0.0622 ± 0.0263 0.1936 ± 0.1503

### v2

@@ I Introduction
  • RQ2: Horizon scaling. How does performance evolve as the forecasting horizon increases (from hours to days ahead)?
  • RQ3: Cross-level generalization. Can TTM generalize effectively to ISP traffic across hierarchical levels (institutions, subnets, and IP addresses), without retraining?
@@
  • Evaluation of TTM on real ISP traffic. We apply IBM’s TTM to CESNET-TimeSeries24 and evaluate its performance under the benchmark protocol. To facilitate reproducibility, we publish the source code of the evaluation on GitHub  [^1] https://github.com/FanL1u/TSFM-for-ISP-traffic-forecasting.git.
@@
  • Negative result on few-shot learning. We systematically evaluate few-shot fine-tuning and exogenous features, and find negligible improvement.
@@ II-A Statistical Forecasting
Early network-traffic forecasting methods relied on statistical time-series models such as exponential smoothing and the ARIMA models [12]. These approaches remain appealing for their interpretability and low computational cost, yet a linear correlation structure is assumed among the time series values, and therefore, no nonlinear patterns can be captured by the ARIMA model. The approximation of linear models to complex real-world problems is not always satisfactory. Many researchers have argued that real-world systems are often nonlinear [31]. Hybrid extensions combining ARIMA with neural models such as ANN and LSTM [30, 16] have been explored to capture both linear and non-linear patterns, but their improvements often come at the expense of increased model complexity and extensive parameter tuning.
@@ II-E Gap in ISP Forecasting
While foundation models have achieved promising results in finance, energy, and transportation [10], their application to Internet traffic remains largely unexplored. Ferreira et al. [9] identify generalization and transfer learning as open challenges in network-traffic prediction.
@@
Our study addresses these questions by systematically evaluating TTM on the CESNET-TimeSeries24 dataset, assessing zero-shot, few-shot, and exogenous-feature settings to show whether pretrained forecasting models can match or surpass specialized deep-learning baselines while maintaining low computational overhead.
@@ III Methodology
This section describes the experimental pipeline used to evaluate TTM-R2 (where R2 denotes the model’s release version; TTM is used throughout the paper) for ISP traffic forecasting. As illustrated in Figure 1, the workflow comprises five stages: (1) dataset preprocessing and window generation, (2) model configuration and context–horizon setup, (3) zero-shot and few-shot execution, (4) metric computation across overlapping windows, and (5) efficiency benchmarking on CPU-only mode. The evaluation workflow is dataset-agnostic and applies to other time-series datasets with fixed sampling intervals and sufficient historical context for context–horizon forecasting.
@@ III-A Dataset and Forecasting Targets
Each record captures multiple traffic attributes including n_flows, n_packets, n_bytes, tcp_udp_ratio_bytes, and avg_duration etc. Data are collected at a base granularity of 10 minutes and further aggregated to hourly and daily resolution by temporal averaging, resulting in three temporal granularities:
@@
Among the available metrics, we focus primarily on n_bytes, representing the total traffic volume per interval. We initially experimented with multiple target attributes and observed consistent performance trends across them. We therefore report n_bytes as the representative target for clarity and comparability with other deep-learning baseline benchmarks [18]. In addition, this metric directly reflects bandwidth utilization and network load, making it operationally relevant for applications such as capacity planning, congestion management, and anomaly detection in ISP networks.
@@ III-C Preprocessing
Input–output tensor format
Each sliding window maps to a single model sample. The input context tensor has shape $[L,1]$ (or $[L,C]$ when exogenous channels are included) and the prediction target has shape $[H,1]$, as illustrated in Figure 2. Among the dataset’s traffic attributes, only the scaled target column is assigned to the input channel; timestamps are used solely for windowing and are not provided to the model. In batched inference the shapes become $B,L,1]\rightarrow[B,H,1]$.
Step Timestamp (hourly) Scaled n_bytes
$t_{1}$ 2024-01-15 00:00 0.312
$t_{2}$ 2024-01-15 01:00 0.287
$\vdots$ $\vdots$ $\vdots$
$t_{L}$ 2024-02-26 07:00 0.445
$t_{L+1}$ 2024-02-26 08:00 0.461 (pred.)
$\vdots$ $\vdots$ $\vdots$
$t_{L+H}$ 2024-03-01 07:00 0.389 (pred.)
Figure 2: Illustration of one sliding window (L=1024L{=}1024, H=96H{=}96, hourly). The top $L$ rows form the context input $[L,1]$; the bottom $H$ rows are the prediction target $[H,1]$. Values shown are after min–max scaling.
@@ III-F Exogenous Variables
We evaluate the effect of exogenous signals by comparing runs with and without binary weekend and holiday flags. The default configuration uses univariate input only.
@@ IV Results and Discussion
IV-A Effect of Context Length (RQ1)
@@
For hourly forecasts ($H=96$), the RMSE remains nearly unchanged as $L$ increases from 512 to 1536, indicating that the model already encodes sufficient temporal memory with moderate context. Similarly, at 10-minute resolution ($H=336$), longer lookback windows show no consistent improvement in either RMSE or $R^{2}$. These results suggest that TTM’s pretrained temporal representations capture recurring diurnal and weekly cycles even when short historical windows are used.
@@
IV-B Effect of Forecast Horizon (RQ2)
@@
TABLE II: Forecasting Accuracy Summary Across All Experimental Configurations. Values reported as mean ± standard deviation after removing top/bottom 5% outliers.
Experiment Context L Horizon H Few-shot % RMSE
Context Length Sensitivity (Hourly, Institutions)
Hourly L=512 512 96 0 0.0566 ± 0.0269 0.2127 ± 0.1388
Hourly L=1024 1024 96 0 0.0551 ± 0.0263 0.2460 ± 0.1571
Hourly L=1536 1536 96 0 0.0551 ± 0.0265 0.2498 ± 0.1576
Cross-Level Generalization (Hourly, L=1024, H=96)
Cross: Institutions 1024 96 0 0.0551 ± 0.0263 0.2460 ± 0.1571
Cross: Subnets 1024 96 0 0.0541 ± 0.0365 0.1685 ± 0.1672
Cross: IPs 1024 96 0 0.0317 ± 0.0262 $1.3872\pm 8.6764$
Horizon Scaling (10-minute, Institutions, L=1024)
10-min H=96 1024 96 0 0.0259 ± 0.0142 0.0975 ± 0.0982
10-min H=336 1024 336 0 0.0268 ± 0.0144 0.0654 ± 0.0725
10-min H=720 1024 720 0 0.0271 ± 0.0146 0.0456 ± 0.0638
Few-Shot Learning (Hourly, Institutions, L=1024, H=96)
Few-shot: 0% (zero-shot) 1024 96 0 0.0551 ± 0.0263 0.2460 ± 0.1571
Few-shot: 10% 1024 96 10 0.0718 ± 0.0380 $0.2170\pm 0.3600$
Few-shot: 30% 1024 96 30 0.0605 ± 0.0299 0.1196 ± 0.1847
Few-shot: 50% 1024 96 50 0.0622 ± 0.0263 0.1936 ± 0.1503
@@
TABLE III: Inference and Training Efficiency Across Experimental Configurations (CPU-only, Apple M2 Pro, 16GB RAM). All values reported as mean ± standard deviation. Training time only applicable to few-shot experiments.
@@
IV-C Cross-Level Generalization (RQ3)
Our findings with TTM align with this general trend but demonstrate greater resilience. As summarized in Table II, mean $R^{2}$ values remain positive and stable for institutions ($\sim$0.25) and subnets ($\sim$0.17), while declining at the IP level ($\sim$–1.39) due to extreme sparsity. Unlike the baseline study, which suggested that distinct modeling strategies may be needed at each aggregation level (e.g., recurrent models for institutions, specialized preprocessing for IPs), TTM maintains a single unified architecture across all levels, without any retraining or architectural modification.
[…]
IV-D Few-Shot Adaptation and Exogenous Variables (RQ4, RQ5)
Figure 6 shows that as the fraction of fine-tuning data increases from 10% to 50% ($L=1024$, $H=96$), RMSE improves modestly (from $\sim$0.075 to $\sim$0.070), while $R^{2}$ remains largely stable. This indicates that limited adaptation can refine magnitude estimation but does not substantially alter the model’s overall calibration.
IV-E Deployability and Efficiency (RQ6)
Computational efficiency is a key requirement for operational ISP deployment. Unlike conventional deep learning models that rely on GPUs and extensive retraining, TTM delivers strong forecasting accuracy entirely on CPU hardware with minimal overhead. All experiments were executed on a MacBook Pro (M2 Pro, 16 GB RAM, CPU-only). Table III summarizes inference and training costs across configurations. For hourly forecasting at the institutional level ($L=1024$, $H=96$), TTM achieves 0.044 s per 100 points ($\sim$2,200 samples/s). Efficiency remains stable across context lengths (L=512L=512–1536) with only a 0.006 s difference.
@@ IV-F Quantitative Summary
Table II reports the aggregated metrics (mean ± std across all evaluated series). TTM maintains consistent accuracy across context lengths and horizons, with RMSE around 0.025–0.056 and stable inference times (Table III) even on CPU-only mode. R² remains positive for institutional and subnet levels, with degradation at the IP level due to flat or zero-variance series. These results confirm the model’s robustness and computational efficiency across traffic hierarchies.
@@ IV-G Limitations and Future Work
Fourth, IP-level performance naturally degrades under extreme sparsity and zero inflation. Several directions could mitigate this limitation in future work. First, temporal aggregation (e.g., daily or weekly rollups) can reduce intermittency at the cost of temporal resolution. Second, two-stage intermittent forecasting, predicting activity occurrence followed by conditional magnitude, provides a principled way to handle zero-inflated traffic series. Third, context augmentation with history-based features (e.g., rolling activity counts, time since last non-zero event) and fine-grained calendar signals may remain informative even during prolonged zero intervals. Finally, hierarchical information sharing, where subnet- or institution-level signals regularize IP-level predictions, could allow fine-grained forecasts to borrow strength from more stable aggregates.
@@ V Conclusion
Overall, our results answer RQ1–RQ6 by showing that TTM is largely insensitive to context length, remains robust as the forecast horizon increases, generalizes across aggregation levels, gains little from few-shot adaptation or simple exogenous features, and delivers practical CPU-only inference efficiency for ISP-scale deployment.